### v1

# On local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms

Abstract.
In this article, we generalize some results in Chan-Yuan (2025) to local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms. In particular, we obtain further rigidity theorems and non-existence theorems for such maps.
2020 Mathematics Subject Classification:
Primary 32H02; Secondary 32Q15, 53C24
The author is partially supported by a program of the Chinese Academy of Sciences
1. Introduction
In Chan-Yuan [CY25] we have studied local holomorphic maps between Kähler manifolds preserving $(p,p)$-forms, namely, the $p$-th exterior power of their Kähler forms. In particular, we have established several rigidity theorems and non-existence theorems. Later on, we realize that by similar methods, some of these results can be generalized to the case of local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms, namely the $p$-th exterior power of their Hermitian forms. In this article, we aim at providing further rigidity theorems and non-existence theorems for local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms.
We will follow the settings in Chan-Yuan [CY25] throughout this article. Recall that a (finite-dimensional) complex space form is a connected complete Kähler manifold of constant holomorphic sectional curvature. We write $X=X(c)$ to indicate that $X$ is of constant holomorphic sectional curvature $c$. The universal covering space of a complex space form of complex dimension $n$ is either the complex Euclidean space $\mathbb{C}^{n}$, the complex unit ball $\mathbb{B}^{n}$ or the complex projective space $\mathbb{P}^{n}$. We say that two complex space forms $M$ and $N$ are of the same type if $M=M(c)$ and $N=N(c^{\prime})$ such that either $c$ and $c^{\prime}$ are of the same sign or $c=c^{\prime}=0$. We say that $M$ and $N$ are of different types if $M$ and $N$ are not of the same type.
In [Ca53], Calabi has proved that there does not exist a local holomorphic isometric embedding between complex space forms of different types. This motivates our study in Chan-Yuan [CY25], and we have partially generalized Calabi’s result to the case of local holomorphic maps preserving $(p,p)$-forms (cf. Chan-Yuan [CY25, Theorem 1.4]). Recently, Arezzo-Li-Loi [ALL25, Corollary 1.7] has proved that there does not exist a Ricci-flat non-compact Kähler submanifold of any finite dim. complex projective space equipped with the Fubini–Study metric. Hence, it follows from Chan-Yuan [CY25, Theorem 1.4 and Remark 3.9] and Arezzo-Li-Loi [ALL25, Corollary 1.7] that if $(M,\omega_{M})$ and $(N,\omega_{N})$ are complex space forms of complex dimensions $m$ and $n$ respectively such that $M$ and $N$ are of different types, then for any real constant $\lambda>0$ and any integer $p$, $1\leq p\leq\min\{m,n\}$, there does not exist a local holomorphic map $F:(M;x_{0})\to(N;F(x_{0}))$ such that $F^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p}$ holds. This yields a complete generalization of Calabi’s result. (Noting that in Chan-Yuan [CY25, Theorem 1.4] we exclude the case where the universal cover of $M$ is biholomorphic to $\mathbb{C}^{m}$, $N=\mathbb{P}^{n}$ and $2\leq p=m<n$.)
2. Preliminaries
Given a Hermitian manifold $M$ of complex dimension $m$, we write $\omega_{M}$ for its Hermitian form, and $g_{M}$ for the corresponding Hermitian metric. Fix $x_{0}\in M$ and let $w=(w_{1},\ldots,w_{m})$ be local holomorphic coordinates around $x_{0}$. For $1\leq p\leq m$, we write ${\bf A}=\{I=(i_{1},\ldots,i_{p}):1\leq i_{1}<\cdots<i_{p}\leq m\}$ and $\omega_{M}^{p}=(\sqrt{-1})^{p}\sum_{I,J\in{\bf A}}(\omega_{M}^{p})_{I\overline{J}}dw^{I}\wedge d\overline{w^{J}}$ in terms of the local holomorphic coordinates. In what follows, for any $v\in\wedge^{p}T_{x_{0}}(M)$, we write $v=\sum_{I\in{\bf A}}v_{I}{\partial\over\partial w_{i_{1}}}\wedge\cdots\wedge{\partial\over\partial w_{i_{p}}}$, $I=(i_{1},\ldots,i_{p})$, and write $\omega_{M}^{p}(x_{0})(v,\overline{v}):=\sum_{I,J\in{\bf A}}(\omega_{M}^{p})_{I\overline{J}}(x_{0})v_{I}\overline{v_{J}}$.
For any complex manifold $X$, we write $T_{X}$ for the holomorphic tangent bundle of $X$, $\Theta_{E}$ for the curvature operator of the Chern connection on a Hermitian holomorphic vector bundle $(E,h)$ over $X$, and $\Lambda(X)$ for the Umehara algebra on $X$ as defined in Umehara [Um88]. More precisely, $\Lambda(X)$ is the associative algebra of real analytic functions on $X$ that consists of real linear combinations of functions $f\overline{g}+g\overline{f}$ for holomorphic functions $f$ and $g$ on $X$.
For any irreducible bounded symmetric domain $D$, we denote by $g_{D}$ the canonical Kähler-Einstein metric on $D$ so that the minimum of the holomorphic sectional curvatures of $(D,g_{D})$ equals 2-2. Given any bounded symmetric domain $\Omega$, we say that $g^{\prime}_{\Omega}$ is a canonical Kähler metric on $\Omega$ whenever $g^{\prime}_{\Omega}=\sum_{j=1}^{k}\lambda_{j}{\rm Pr}_{j}^{*}g_{\Omega_{j}}$ for some real constants $\lambda_{j}>0$, $1\leq j\leq k$, where $\Omega_{j}$, $1\leq j\leq k$, are the irreducible factors of $\Omega$, and ${\rm Pr}_{j}:\Omega\to\Omega_{j}$ is the canonical projection onto the $j$-th irreducible factor $\Omega_{j}$, $1\leq j\leq k$. We denote by $g_{\mathbb{P}^{n}}$ the Fubini-Study metric on the complex projective space $\mathbb{P}^{n}$ of complex dimension $n$ such that $(\mathbb{P}^{n},g_{\mathbb{P}^{n}})$ is of constant holomorphic sectional curvature $2$.
3. Rigidity theorem for local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms
Let $(X,g)$ be a complex $n$-dimensional Hermitian manifold with the Hermitian form $\omega$, namely, $\omega=g(J\cdot,\cdot)$ with $J$ being the induced almost complex structure on $(X,g)$. For simplicity, in the present article we write $(X,\omega)$ to indicate that $\omega$ is the Hermitian form. In terms of local holomorphic coordinates $(z_{1},\ldots,z_{n})$ on $X$, we may write $\omega=\sqrt{-1}\sum_{i,j=1}^{n}g_{i\overline{j}}dz_{i}\wedge d\overline{z_{j}}$ locally. Then, for $1\leq p\leq n$ we have
$\omega^{p}=(\sqrt{-1})^{p}p!\sum_{\tiny\begin{split}&1\leq i_{1}<\cdots<i_{p}\leq n,\\ &1\leq j_{1}<\cdots<j_{p}\leq n\end{split}}\det\begin{pmatrix}g_{i_{s}\overline{j_{t}}}\end{pmatrix}_{1\leq s,t\leq p}(dz_{i_{1}}\wedge d\overline{z_{j_{1}}})\wedge\cdots\wedge(dz_{i_{p}}\wedge d\overline{z_{j_{p}}}).$
Let $(M,g_{M})$ and $(N,g_{N})$ be Hermitian manifolds of complex dimensions $m$ and $n$ respectively. Fix some point $x_{0}\in M$. Let $F:(M;x_{0})\to(N;F(x_{0}))$ be a germ of holomorphic map such that
(3.1) $F^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p},$
for some integer $p$, $1\leq p\leq\min\{m,n\}$, and some real constant $\lambda>0$. Then, such a germ of holomorphic map $F$ is of rank at least $p$ near $x_{0}$.
Write $w=(w_{1},\ldots,w_{m})$ (resp. $z=(z_{1},\ldots,z_{n})$) for the local holomorphic coordinates on $M$ (resp. $N$) around the point $x_{0}$ (resp. $F(x_{0})$). We write
$\omega_{M}=\sqrt{-1}\sum_{i,j=1}^{m}(g_{M})_{i\overline{j}}dw_{i}\wedge d\overline{w_{j}},\quad\omega_{N}=\sqrt{-1}\sum_{i,j=1}^{n}(g_{N})_{i\overline{j}}dz_{i}\wedge d\overline{z_{j}},$
$F(w)=(F_{1}(w),\ldots,F_{n}(w))$ locally around $x_{0}$ in terms of the local holomorphic coordinates, and
${\partial(F_{i_{1}},\ldots,F_{i_{p}})\over\partial(w_{l_{1}},\ldots,w_{l_{p}})}:=\begin{pmatrix}{\partial F_{i_{1}}\over\partial w_{l_{1}}}&\cdots&{\partial F_{i_{1}}\over\partial w_{l_{p}}}\\ \vdots&\ddots&\vdots\\ {\partial F_{i_{p}}\over\partial w_{l_{1}}}&\cdots&{\partial F_{i_{p}}\over\partial w_{l_{p}}}\end{pmatrix}$
for $1\leq i_{1}<\cdots<i_{p}\leq n$ and $1\leq l_{1}<\cdots<l_{p}\leq m$. Then, we have
$F^{*}\omega_{N}^{p}=C_{p}\cdot\sum_{\tiny\begin{split}&1\leq l_{1}<\cdots<l_{p}\leq m,\\ &1\leq k_{1}<\cdots<k_{p}\leq m\end{split}}\Theta_{l_{1},\ldots,l_{p};\overline{k_{1}},\ldots,\overline{k_{p}}}(w)dw_{l_{1}}\wedge\cdots\wedge dw_{l_{p}}\wedge d\overline{w_{k_{1}}}\wedge\cdots\wedge d\overline{w_{k_{p}}},$
where $C_{p}:=(\sqrt{-1})^{p}(-1)^{p(p-1)\over 2}p!$ and
$\Theta_{l_{1},\ldots,l_{p};\overline{k_{1}},\ldots,\overline{k_{p}}}(w):=\sum_{\tiny\begin{split}&1\leq i_{1}<\cdots<i_{p}\leq n,\\ &1\leq j_{1}<\cdots<j_{p}\leq n\end{split}}\det\begin{pmatrix}(g_{N})_{i_{s}\overline{j_{t}}}(F(w))\end{pmatrix}_{1\leq s,t\leq p}\hskip 56.9055pt$
$\hskip 142.26378pt\cdot\det\left({\partial(F_{i_{1}},\ldots,F_{i_{p}})\over\partial(w_{l_{1}},\ldots,w_{l_{p}})}\right)\det\left(\overline{{\partial(F_{j_{1}},\ldots,F_{j_{p}})\over\partial(w_{k_{1}},\ldots,w_{k_{p}})}}\right).$
In terms of the local holomorphic coordinates, (3.1) becomes
(3.2) $\Theta_{l_{1},\ldots,l_{p};\overline{k_{1}},\ldots,\overline{k_{p}}}(w)=\lambda\det\begin{pmatrix}(g_{M})_{l_{s}\overline{k_{t}}}(w)\end{pmatrix}_{1\leq s,t\leq p}$
for $1\leq l_{1}<\cdots<l_{p}\leq m$ and $1\leq k_{1}<\cdots<k_{p}\leq m$.
By the same proof of Proposition 3.1 in Chan-Yuan [CY25], we have the following rigidity result on local holomorphic maps from $M$ to $N$ between Hermitian manifolds $M$ and $N$ preserving $(p,p)$-forms when $1\leq p<\dim_{\mathbb{C}}(M)$.
Proposition 3.1 (cf. Chan-Yuan [CY25, Proposition 3.1]).
Let $(M,\omega_{M})$ and ((NN, $\omega_{N}$)) be Hermitian manifolds of finite complex dimensions $m$ and $n$ respectively, where $\omega_{M}$ and $\omega_{N}$ denote the corresponding Hermitian forms. Let $F:(M;x_{0})\to(N;F(x_{0}))$ be a germ of holomorphic map such that $F^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p}$ for some real constant $\lambda>0$ and some integer $p$, $1\leq p\leq\min\{m,n\}$. Then, $m\leq n$.
If we assume in addition that $1\leq p<m$, then $F^{*}\omega_{N}=\lambda^{1\over p}\omega_{M}$ so that $F:(M,\lambda^{1\over p}\omega_{M};x_{0})\to(N,\omega_{N};F(x_{0}))$ is a local holomorphic isometry.
4. Non-existence Theorems
4.1. The unit sphere bundles on Hermitian manifolds
Let $(M,\omega_{M})$ and $(N,\omega_{N})$ be Hermitian manifolds of complex dimensions $m$ and $n$ respectively. Let $F:(M;x_{0})\to(N;F(x_{0}))$ be a germ of holomorphic map such that $F^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p}$ for some integer $p$, $1\leq p\leq\min\{m,n\}$, and some real constant $\lambda>0$. Write $U$ for an open neighborhood of $x_{0}$ in the domain of $F$ in which we have the holomorphic coordinates $(w_{1},\ldots,w_{m})$ on $U$. Define the unit sphere bundles
(4.1) $S_{1}:=\left\{(w,v)\in\wedge^{p}T_{U}:\lambda\omega_{M}^{p}(w)(v,\overline{v})=1\right\},\;S_{2}:=\left\{(z,\xi)\in\wedge^{p}T_{N}:\omega_{N}^{p}(z)(\xi,\overline{\xi})=1\right\}.$
The map $F$ induces the holomorphic map $f:=(F,dF):\wedge^{p}T_{U}\to\wedge^{p}T_{N}$ defined by $f(w,v):=(F(w),dF_{w}(v))$ for all $(w,v)\in\wedge^{p}T_{U}$. By (3.1), $f$ maps $S_{1}$ into $S_{2}$. Writing $\rho_{1}(w,v):=\lambda\omega_{M}^{p}(w)(v,\overline{v})-1$ and $\rho_{2}(z,\xi):=\omega_{N}^{p}(z)(\xi,\overline{\xi})-1$, we define
$U_{1}:=\left\{(w,v)\in\wedge^{p}T_{U}:\rho_{1}(w,v)<0\right\},\;V_{1}:=\left\{(w,v)\in\wedge^{p}T_{U}:\rho_{1}(w,v)>0\right\}$
and
$U_{2}:=\left\{(z,\xi)\in\wedge^{p}T_{N}:\rho_{2}(z,\xi)<0\right\},\;V_{2}:=\left\{(z,\xi)\in\wedge^{p}T_{N}:\rho_{2}(z,\xi)>0\right\}.$
Then, by (3.1) and the definitions of $\rho_{1}$ and $\rho_{2}$, we have $\rho_{2}\circ f=\rho_{1}$ so that $f=(F,dF)$ maps $U_{1}$ into $U_{2}$, and maps $V_{1}$ into $V_{2}$.
As in Chan-Yuan [CY25], we make use of the method of CR geometry to study the CR map $f:S_{1}\to S_{2}$. We have the following non-existence theorem by arguments similar to those in Chan-Yuan [CY25, Proof of Theorem 3.2].
Theorem 4.1.
Let $(M,\omega_{M})$ and $(N,\omega_{N})$ be Hermitian manifolds of complex dimensions $m$ and $n$ respectively such that $(K_{M}^{*},\det(g_{M}))$ is a positive Hermitian holomorphic line bundle, where $K_{X}^{*}$ denotes the anti-canonical line bundle of any complex manifold $X$ and $g_{M}$ denotes the Hermitian metric of $(M,\omega_{M})$. Then, we have the following.
  1. (a)
    If $(N,\omega_{N})$ is of semi-negative holomorphic bisectional curvature, then there does not exist any germ of holomorphic map $F:(M;x_{0})\to(N;F(x_{0}))$ such that $F^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p}$ holds, where $\lambda>0$ is a real constant and $p$ is an integer with $1\leq p\leq\min\{m,n\}$.
  2. (b)
    If $(N,\omega_{N})$ is a period domain equipped with the canonical invariant metric, then there does not exist any germ of horizontal holomorphic map $F:(M;x_{0})\to(N;F(x_{0}))$ such that $F^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p}$ holds, where $\lambda>0$ is a real constant and $p$ is an integer with $1\leq p\leq\min\{m,n\}$.
Proof.
Assume the contrary that there exists a germ of holomorphic map $F:(M;x_{0})\to(N;F(x_{0}))$ such that $F^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p}$ holds. Write $U\subset M$ for the domain of $F$. Then, in the above settings we have the CR map $f:=(F,dF):S_{1}\to S_{2}$ between the unit sphere bundles $S_{1}$ and $S_{2}$ defined in (4.1), and we write $\rho_{1}$ and $\rho_{2}$ for the defining functions of $S_{1}$ and $S_{2}$ respectively as in the above. We have $\rho_{2}\circ f=\rho_{1}$, by $F^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p}$. Fix a point $(w_{0},v_{0})\in S_{1}$. By Chan-Yuan [CY25, Lemma 2.4], for any $\eta,\eta^{\prime}\in T^{1,0}_{(w_{0},v_{0})}(S_{1})$, we have
(4.2) $(\partial\overline{\partial}\rho_{2})(f(w_{0},v_{0}))(f_{*}(\eta),\overline{f_{*}(\eta^{\prime})})=(\partial\overline{\partial}\rho_{1})(w_{0},v_{0})(\eta,\overline{\eta^{\prime}}).$
We consider Case (a). By computations similar to those in Chan-Yuan [CY25, Section 2.3] and Yuan [Yu17], since $(N,\omega_{N})$ is of semi-negative holomorphic bisectional curvature, $S_{2}$ is weakly pseudoconvex. By the identity $^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p}, FF$ is of rank $\geq p$ at every point in $U$. Thus, we may choose a vector $\eta=(\eta_{1},{\bf 0})\in T^{1,0}_{(w_{0},v_{0})}(S_{1})$ with $\eta_{1}\in T_{w_{0}}(U)$ and $f_{*}(\eta)=(dF_{w_{0}}(\eta_{1}),\eta^{\prime})\neq{\bf 0}$. If $1\leq p<m$, then by Proposition 3.1 $F$ would be a germ of holomorphic isometry, i.e., $F^{*}\omega_{N}=\lambda^{1\over p}\omega_{M}$, and thus $F^{*}\omega_{N}^{m}=\lambda^{m\over p}\omega_{M}^{m}$. Therefore, we may assume without loss of generality that $p=m$. Then, $F:U\to N$ is a holomorphic immersion. By the assumption on the anti-canonical line bundle $K_{M}^{*}$, $(K_{U}^{*},\det(g_{M})|_{K_{U}^{*}})$ is a positive Hermitian holomorphic line bundle over $U$, and thus by computations similar to those in Chan-Yuan [CY25] we have
$(\partial\overline{\partial}\rho_{1})(w_{0},v_{0})(\eta,\overline{\eta})=-C\Theta_{K_{U}^{*}}(\eta_{1},\overline{\eta_{1}},v_{0},\overline{v_{0}})<0$
for some real constant $C>0$. On the other hand, $(\partial\overline{\partial}\rho_{2})(f(w_{0},v_{0}))(f_{*}(\eta),\overline{f_{*}(\eta)})\geq 0$ since $S_{2}$ is weakly pseudoconvex and $f_{*}(\eta)\neq{\bf 0}$. This clearly contradicts with (4.2), and the proof of the first assertion is complete.
Now, we consider Case (b), i.e., $(N,\omega_{N})$ is a period domain equipped with the canonical invariant metric and $F$ is a horizontal holomorphic map, i.e., $F:U\to N$ is a holomorphic map such that $dF_{x}(T_{x}(M))$ lies inside the horizontal tangent space $T^{h}_{F(x)}(N)\subset T_{F(x)}(N)$ for all $x\in U$. Then, $F(U)\subset N$ is an integral manifold for the horizontal distribution on the period domain $N$. By a result of Carlson-Toledo [CT88, Proposition 5.3], we have $dF_{x}(T_{x}(M)),dF_{x}(T_{x}(M))]=0$ for $x\in U$. It follows from Peters [Pet90, Corollary 1.8] that the holomorphic bisectional curvature $R_{v\bar{v}\xi\bar{\xi}}(N,\omega_{N})\leq 0$ along the directions $v,\xi\in dF_{x}(T_{x}(M))$, $x\in U$. Hence, we still have $(\partial\overline{\partial}\rho_{2})(f(w_{0},v_{0}))(f_{*}(\eta),\overline{f_{*}(\eta)})$ $\geq$ 0. By the same argument as in the above, this contradicts with (4.2), and the proof of the second assertion is complete. ∎
Corollary 4.2.
Let $(M,\omega_{M})$ be a Hermitian manifold such that $(K_{M}^{*},\det(g_{M}))$ is a positive Hermitian holomorphic line bundle. Write $m:=\dim_{\mathbb{C}}(M)$. Let $\Omega\Subset\mathbb{C}^{n}$ be a bounded symmetric domain with $\omega_{\Omega}$ being the Kähler form of the canonical Kähler metric $\sum_{j=1}^{k}\lambda_{j}{\rm Pr}_{j}^{*}g_{\Omega_{j}}$, where $\Omega_{j}$, $1\leq j\leq k$, are the irreducible factors of $\Omega$, ${\rm Pr}_{j}:\Omega\to\Omega_{j}$ is the canonical projection onto the $j$-th irreducible factor $\Omega_{j}$, and $\lambda_{j}>0$, $1\leq j\leq k$. Then, there does not exist any germ of holomorphic map $F:(M;x_{0})\to(\Omega;F(x_{0}))$ such that $F^{*}\omega_{\Omega}^{p}=\lambda\omega_{M}^{p}$ holds, where $\lambda>0$ is a real constant and $p$ is an integer with $1\leq p\leq\min\{m,n\}$.
Proof.
It is clear that $(\Omega,\omega_{\Omega})$ is of semi-negative holomorphic bisectional curvature, and thus the result follows from Theorem 4.1. ∎
4.2. The Umehara algebra on bounded symmetric domains and its applications
We have the following basic fact on Umehara algebra.
Lemma 4.3.
Let $D\Subset\mathbb{C}^{m}$ be a bounded symmetric domain equipped with the canonical Kähler metric $g^{\prime}_{D}:=\sum_{j=1}^{k}\lambda_{j}{\rm Pr}_{j}^{*}g_{D_{j}}$ for some $\lambda_{j}>0$, $1\leq j\leq k$, where $D_{j}$, $1\leq j\leq k$ are the irreducible factors of $D$, and ${\rm Pr}_{j}:D\to D_{j}$ is the canonical projection onto the $j$-th irreducible factor $D_{j}$, $1\leq j\leq k$. Write $\omega_{D}$ for the Kähler form of $(D,g^{\prime}_{D})$ and $m_{j}:=\dim_{\mathbb{C}}(D_{j})$. The real-analytic function
$\phi_{m}(w):=\det\left((g^{\prime}_{D})_{s\bar{t}}(w)\right)_{1\leq s,t\leq m}$
on $D$ does not lie inside the Umehara algebra $\Lambda(D)$.
Proof.
Assume the contrary that $\phi_{m}\in\Lambda(D)$. We choose the Harish-Chandra coordinates $($$w_{1}$$\ldots,$ $w_{m}$)) $\in$ $\mathbb{C}^{m}$ on the bounded symmetric domain $D$ so that $D\cap\{(w_{1},\ldots,w_{m})\in\mathbb{C}^{m}:w_{j}=0\text{ for }j=2,\ldots,m\}$ is the minimal disk $\Delta_{0}$. By restricting to $\Delta_{0}$ we obtain a function $\widetilde{\phi}_{m}(\zeta):=\phi_{m}(\zeta,0,\ldots,0)$ for $\zeta\in\{w_{1}\in\mathbb{C}:(w_{1},0\ldots,0)\in D\}=\{w_{1}\in\mathbb{C}:|w_{1}|^{2}<1\}=:\Delta$. Then, we would have $\widetilde{\phi}_{m}\in\Lambda(\Delta)$ by the assumption. We put $(z^{1},\ldots,z^{k})=(w_{1},\ldots,w_{m})$ with $z^{j}=(z^{j}_{1},\ldots,z^{j}_{m_{j}})$ for $1\leq j\leq k$. We have
$\omega_{D}=-\sqrt{-1}\sum_{j=1}^{k}\lambda_{j}\partial\overline{\partial}\log\rho_{j}$
where
$\rho_{j}=1-\sum_{\mu=1}^{m_{j}}|z^{j}_{\mu}|^{2}+\sum_{l=1}^{N^{\prime}_{j}}(-1)^{\deg G^{j}_{l}}|G^{j}_{l}(z^{j})|^{2}$
for some homogeneous polynomials $G^{j}_{l}$, $1\leq l\leq N^{\prime}_{j}$, of degree $\geq 2$ (cf. Chan-Xiao-Yuan [CXY17, Section 4]). We compute
$\widetilde{\phi}_{m}(\zeta)=\det\begin{bmatrix}{\lambda_{1}\over(1-|\zeta|^{2})^{2}}&&&&\\ &{\lambda_{1}\over 1-|\zeta|^{2}}{\bf I}_{m_{1}-1}&&&\\ &&\lambda_{2}{\bf I}_{m_{2}}&&\\ &&&\ddots&\\ &&&&\lambda_{k}{\bf I}_{m_{k}}\end{bmatrix}={\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{m_{1}+1}},$
which does not lie in $\Lambda(\Delta)$ by [Um88], a plain contradiction. Hence, $\phi_{m}\not\in\Lambda(D)$. ∎
By using Lemma 4.3, we have the following non-existence result by using the Umehara algebra.
Proposition 4.4.
Let $D\Subset\mathbb{C}^{m}$ be a bounded symmetric domain equipped with the canonical Kähler metric $g^{\prime}_{D}:=\sum_{j=1}^{k}\lambda_{j}{\rm Pr}_{j}^{*}g_{D_{j}}$ for some $\lambda_{j}>0$, $1\leq j\leq k$, where $D_{j}$, $1\leq j\leq k$ are the irreducible factors of $D$, and ${\rm Pr}_{j}:D\to D_{j}$ is the canonical projection onto the $j$-th irreducible factor $D_{j}$, $1\leq j\leq k$. Write $\omega_{D}$ for the Kähler form of $(D,g^{\prime}_{D})$ and $m_{j}:=\dim_{\mathbb{C}}(D_{j})$. Let $X$ be a projective manifold of complex dimension $n$ equipped with the Kähler metric $g_{X}$ such that there exists a holomorphic isometric embedding $(X,g_{X})\hookrightarrow(\mathbb{P}^{N_{1}},\mu_{1}g_{\mathbb{P}^{N_{1}}})\times\cdots\times(\mathbb{P}^{N_{l}},\mu_{l}g_{\mathbb{P}^{N_{l}}})$ for some real constants $\mu_{j}>0$, $1\leq j\leq l$, and some integers $N_{j}\geq 1$, $1\leq j\leq l$, where $l$ is a positive integer. Then, for any $p$, $1\leq p\leq\min\{m,n\}$, there does not exist a germ of holomorphic map $F:(D;x_{0})\to(X;F(x_{0}))$ such that
(4.3) $F^{*}\omega_{X}^{p}=\lambda\omega_{D}^{p}$
for some real constant $\lambda>0$.
Proof.
Assume the contrary that there exists such a germ of holomorphic map $F$ such that (4.3) holds. The existence of such a map $F$ would imply the existence of a germ of holomorphic map $\hat{F}$ from $D$ to $Y:=\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{l}}$ such that $\hat{F}^{*}\omega_{Y}^{p}=\lambda\omega_{D}^{p}$, where $\omega_{Y}$ denotes the Kähler form of $(\mathbb{P}^{N_{1}},\mu_{1}g_{\mathbb{P}^{N_{1}}})\times\cdots\times(\mathbb{P}^{N_{l}},\mu_{l}g_{\mathbb{P}^{N_{l}}})$. Thus, from now on we may assume $(X,g_{X})=(\mathbb{P}^{N_{1}},\mu_{1}g_{\mathbb{P}^{N_{1}}})\times\cdots\times(\mathbb{P}^{N_{l}},\mu_{l}g_{\mathbb{P}^{N_{l}}})$ and identify $X=Y=\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{l}}$. By Proposition 3.1, if $1\leq p<m$, then $F$ would be a local holomorphic isometry from $(D,\lambda^{1\over p}g^{\prime}_{D})$ to $(X,g_{X})=(\mathbb{P}^{N_{1}},\mu_{1}g_{\mathbb{P}^{N_{1}}})\times\cdots\times(\mathbb{P}^{N_{l}},\mu_{l}g_{\mathbb{P}^{N_{l}}})$, i.e., $F^{*}\omega_{X}=\lambda^{1\over p}\omega_{D}$, and thus we may take the $m$-th exterior power to get $F^{*}\omega_{X}^{m}=\lambda^{m\over p}\omega_{D}^{m}$. Therefore, we may assume without loss of generality that $p=m=\dim_{\mathbb{C}}(D)$.
Let $U\subset D$ be the domain of $F$. Write $F=(F^{1},\ldots,F^{l})$ so that $F^{j}$ maps $U$ into $\mathbb{P}^{N_{j}}$ for $1\leq j\leq l$. We may assume $x_{0}={\bf 0}$ and $F^{j}(x_{0})=[1,0,\ldots,0]$ for $1\leq j\leq l$ without loss of generality by composing $F$ with holomorphic isometries of $(D,g^{\prime}_{D})$ and $(\mathbb{P}^{N_{1}},\mu_{1}g_{\mathbb{P}^{N_{1}}})\times\cdots\times(\mathbb{P}^{N_{l}},\mu_{l}g_{\mathbb{P}^{N_{l}}})$. We also write $N:=\sum_{j=1}^{l}N_{j}$ and $F=(F_{1},\ldots,F_{N})$ for convention. Now, (4.3) is equivalent to
(4.4) $\begin{split}&\sum_{I,J\in{\bf A}}\det((g_{X})_{i_{s}\overline{j_{t}}}(F(w)))_{1\leq s,t\leq p}\det\left({\partial(F_{i_{1}},\ldots,F_{i_{p}})\over\partial(w_{l_{1}},\ldots,w_{l_{p}})}\right)\det\left(\overline{{\partial(F_{j_{1}},\ldots,F_{j_{p}})\over\partial(w_{k_{1}},\ldots,w_{k_{p}})}}\right)\\ =&\lambda\det\left((g^{\prime}_{D})_{l_{s}\overline{k_{t}}}(w)\right)_{1\leq s,t\leq p}\end{split}$
for $1\leq l_{1}<\cdots<l_{p}\leq m$ and $1\leq k_{1}<\cdots<k_{p}\leq m$. Note that
$\begin{pmatrix}(g_{X})_{i\overline{j}}(F(w))\end{pmatrix}_{1\leq i,j\leq N}=\begin{bmatrix}M_{1}(w)&&\\ &\ddots&\\ &&M_{l}(w)\end{bmatrix}$,
where $M_{j}(w):=\mu_{j}\begin{pmatrix}(g_{\mathbb{P}^{N_{j}}})_{s\overline{t}}(F^{j}(w))\end{pmatrix}_{1\leq s,t\leq N_{j}}$ for $1\leq j\leq l$. By similar observation in Chan-Yuan [CY25, Proof of Proposition 3.7] and (4.4), we have
$\left(\prod_{j=1}^{l}(1+\lVert F^{j}(w)\rVert^{2})^{2p}\right)\cdot\det\left((g^{\prime}_{D})_{l_{s}\overline{k_{t}}}(w)\right)_{1\leq s,t\leq p}\in\Lambda(U^{\prime})$
for $1\leq l_{1}<\cdots<l_{p}\leq m$ and $1\leq k_{1}<\cdots<k_{p}\leq m$, where $U^{\prime}$ is a simply connected open neighborhood of $\bf 0$} in $U$ and $\lVert F^{j}(w)\rVert$ denotes the standard complex Euclidean norm of $F^{j}(w)$ in $\mathbb{C}^{N_{j}}$. Let $U^{\prime\prime}$ be a simply connected open neighborhood of $\bf 0$} in $\{\zeta\in\mathbb{C}:(\zeta,{\bf 0})\in U^{\prime}\}$. Now, we put $p=m$. By restricting to the minimal disk $\Delta_{0}:=D\cap\{(w_{1},\ldots,w_{m})\in\mathbb{C}^{m}:w_{j}=0\text{ for }j=2,\ldots,m\}$, we have $\phi_{m}(\zeta,{\bf 0})={\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{m_{1}+1}}$ from the proof of Lemma 4.3, and thus by Chan-Yuan [CY25, Proof of Proposition 3.7] we have
$\psi_{m}(\zeta):={\prod_{j=1}^{l}(1+\lVert F^{j}(\zeta,{\bf 0})\rVert^{2})^{2m}\over(1-|\zeta|^{2})^{m_{1}+1}}\in\Lambda(U^{\prime\prime}).$
It remains to show that $\psi_{m}\not\in\Lambda(U^{\prime\prime})$ in order to prove the non-existence of such a germ of holomorphic map $F$. Actually, by the same arguments in Chan-Yuan [CY25, Proof of Proposition 3.7, p. 635], $\psi_{m}(\zeta)$ would be of infinite rank so that $\psi_{m}\not\in\Lambda(U^{\prime\prime})$. This leads to a contradiction, and thus we conclude that there does not exist such a germ of holomorphic map $F$. ∎
Remark 4.5.
One may ask if this non-existence theorem holds for a rational homogeneous manifold $X\cong G/P$ if we replace the restricted metric $g_{X}$ on $X\subset\mathbb{P}^{N}$ by the $G$-invariant canonical metric $h$ on $X$. The issue is that we do not have much information on the Kähler form of $(X,h)$ for checking things related to the Umehara algebra.
4.3. Local holomorphic maps between Hermitian symmetric manifolds of the compact type and bounded symmetric domains
Let $X$ be a Hermitian symmetric manifold of the compact type. Write $X=X_{1}\times\cdots\times X_{l}$, where each $(X_{j},g_{j})$, $1\leq j\leq l$, is an irreducible Hermitian symmetric manifold of the compact type whose maximum holomorphic sectional curvature equals $2$. A Kähler metric $g_{X}$ on $X$ is said to be a canonical Kähler metric if $g_{X}=\sum_{j=1}^{l}\mu_{j}{\rm Pr}_{j}^{*}g_{j}$ for some real constants $\mu_{j}>0$, $1\leq j\leq l$, where ${\rm Pr}_{j}:X\to X_{j}$ denotes the canonical projection onto the $j$-th irreducible factor $X_{j}$. We fix such a canonical Kähler metric $g_{X}$ on $X$. By the Nakagawa-Takagi theorem (cf. Mok [Mo89, Theorem 1, p. 135]), there exists a holomorphic isometric embedding $\iota_{j}:(X_{j},g_{j})\hookrightarrow(\mathbb{P}^{N_{j}},g_{\mathbb{P}^{N_{j}}})$ for $1\leq j\leq l$. Hence, we have a holomorphic isometric embedding $\iota:(X,g_{X})\hookrightarrow(\mathbb{P}^{N_{1}},\mu_{1}g_{\mathbb{P}^{N_{1}}})\times\cdots\times(\mathbb{P}^{N_{l}},\mu_{l}g_{\mathbb{P}^{N_{l}}})$ given by
$\iota(x_{1},\ldots,x_{l}):=(\iota_{1}(x_{1}),\ldots,\iota_{l}(x_{l}))\quad\forall\;x_{j}\in X_{j},1\leq j\leq l.$
By Corollary 4.2 and Proposition 4.4, we have the following non-existence theorem of local holomorphic maps between Hermitian symmetric manifolds of the compact type and bounded symmetric domains preserving $(p,p)$-forms.
Proposition 4.6.
Let $(X,\omega_{X})$ be a Hermitian symmetric manifold of the compact type equipped with the canonical Kähler metric $g_{X}$. Write $m:=\dim_{\mathbb{C}}(X)$. Let $\Omega\Subset\mathbb{C}^{n}$ be a bounded symmetric domain and let $\omega_{\Omega}$ be the Kähler form of a canonical Kähler metric on $\Omega$. If either (i) $M=\Omega$ and $N=X$ or (ii) $M=X$ and $N=\Omega$, then there does not exist any germ of holomorphic map $F:(M;x_{0})\to(N;F(x_{0}))$ such that $F^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p}$ holds, where $\lambda>0$ is a real constant and $p$ is an integer with $1\leq p\leq\min\{m,n\}$.
Proof.
Since $(X,g_{X})$ is of semi-positive holomorphic bisectional curvature and of positive holomorphic sectional curvature, $(K_{X}^{*},\det(g_{X}))$ is a positive Hermitian holomorphic line bundle over $X$. In addition, there is a holomorphic isometric embedding $\iota:(X,g_{X})\hookrightarrow(\mathbb{P}^{N_{1}},\mu_{1}g_{\mathbb{P}^{N_{1}}})\times\cdots\times(\mathbb{P}^{N_{l}},\mu_{l}g_{\mathbb{P}^{N_{l}}})$ from the construction. Hence, the result follows directly from Corollary 4.2 and Proposition 4.4. ∎
Acknowledgment
We would like to thank Dr. Kwok-Kin Wong for helpful discussions, and Prof. Yuan Yuan for valuable comments. We would also like to thank the referee for helpful comments.

### v2

@@
 Abstract.
-In this article, we generalize some results in Chan-Yuan (2025) to local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms.
+In this article, we generalize some results in Chan-Yuan [Ann.
+Sc.
+Norm.
+Super.
+Pisa Cl.
+Sci.
+(5) 26 (2025), 619–644] to local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms.
 In particular, we obtain further rigidity theorems and non-existence theorems for such maps.
@@
 Chan-Yuan [CY25, Theorem 1.4]).
-Recently, Arezzo-Li-Loi [ALL25, Corollary 1.7] has proved that there does not exist a Ricci-flat non-compact Kähler submanifold of any finite dim.
+Recently, Arezzo-Li-Loi [ALL25, Corollary 1.7] have proved that there does not exist a Ricci-flat non-compact Kähler submanifold of any finite dim.
 complex projective space equipped with the Fubini–Study metric.
@@
 $\rho_{j}=1-\sum_{\mu=1}^{m_{j}}|z^{j}_{\mu}|^{2}+\sum_{l=1}^{N^{\prime}_{j}}(-1)^{\deg G^{j}_{l}}|G^{j}_{l}(z^{j})|^{2}$
-for some homogeneous polynomials $G^{j}_{l}$, $1\leq l\leq N^{\prime}_{j}$, of degree $\geq 2$ (cf. Chan-Xiao-Yuan [CXY17, Section 4]). We compute
-$\widetilde{\phi}_{m}(\zeta)=\det\begin{bmatrix}{\lambda_{1}\over(1-|\zeta|^{2})^{2}}&&&&\\ &{\lambda_{1}\over 1-|\zeta|^{2}}{\bf I}_{m_{1}-1}&&&\\ &&\lambda_{2}{\bf I}_{m_{2}}&&\\ &&&\ddots&\\ &&&&\lambda_{k}{\bf I}_{m_{k}}\end{bmatrix}={\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{m_{1}+1}},$
-which does not lie in $\Lambda(\Delta)$ by [Um88], a plain contradiction. Hence, $\phi_{m}\not\in\Lambda(D)$. ∎
+for some homogeneous polynomials $G^{j}_{l}$, $1\leq l\leq N^{\prime}_{j}$, of degree $\geq 2$ (cf. Chan-Xiao-Yuan [CXY17, Section 4]). By using Mok [Mo14, Lemma 3 and Lemma 5], we compute
+$\begin{split}\widetilde{\phi}_{m}(\zeta)=&\det\begin{bmatrix}{\lambda_{1}\over(1-|\zeta|^{2})^{2}}&&&&&\\ &{\lambda_{1}\over 1-|\zeta|^{2}}{\bf I}_{p_{1}}&&&&\\ &&\lambda_{1}{\bf I}_{m_{1}-p_{1}-1}&&&\\ &&&\lambda_{2}{\bf I}_{m_{2}}&&\\ &&&&\ddots&\\ &&&&&\lambda_{k}{\bf I}_{m_{k}}\end{bmatrix}\\ =&{\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{p_{1}+2}},\end{split}$
+which does not lie in $\Lambda(\Delta)$ by Umehara [Um88], a plain contradiction. Here, $p_{1}$ is a non-negative integer depending only on $D_{1}$. Hence, $\phi_{m}\not\in\Lambda(D)$. ∎
 By using Lemma 4.3, we have the following non-existence result by using the Umehara algebra.
@@
 Now, we put $p=m$.
-By restricting to the minimal disk $\Delta_{0}:=D\cap\{(w_{1},\ldots,w_{m})\in\mathbb{C}^{m}:w_{j}=0\text{ for }j=2,\ldots,m\}$, we have $\phi_{m}(\zeta,{\bf 0})={\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{m_{1}+1}}$ from the proof of Lemma 4.3, and thus by Chan-Yuan [CY25, Proof of Proposition 3.7] we have
-$\psi_{m}(\zeta):={\prod_{j=1}^{l}(1+\lVert F^{j}(\zeta,{\bf 0})\rVert^{2})^{2m}\over(1-|\zeta|^{2})^{m_{1}+1}}\in\Lambda(U^{\prime\prime}).$
+By restricting to the minimal disk $\Delta_{0}:=D\cap\{(w_{1},\ldots,w_{m})\in\mathbb{C}^{m}:w_{j}=0\text{ for }j=2,\ldots,m\}$, we have $\phi_{m}(\zeta,{\bf 0})={\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{p_{1}+2}}$ from the proof of Lemma 4.3, and thus by Chan-Yuan [CY25, Proof of Proposition 3.7] we have
+$\psi_{m}(\zeta):={\prod_{j=1}^{l}(1+\lVert F^{j}(\zeta,{\bf 0})\rVert^{2})^{2m}\over(1-|\zeta|^{2})^{p_{1}+2}}\in\Lambda(U^{\prime\prime}).$
 It remains to show that $\psi_{m}\not\in\Lambda(U^{\prime\prime})$ in order to prove the non-existence of such a germ of holomorphic map $F$.

### v3

@@
 Let $(X,g)$ be a complex $n$-dimensional Hermitian manifold with the Hermitian form $\omega$, namely, $\omega=g(J\cdot,\cdot)$ with $J$ being the induced almost complex structure on $(X,g)$.
-For simplicity, in the present article we write $(X,\omega)$ to indicate that $\omega$ is the Hermitian form.
+For simplicity, in the present article we write $(X,\omega)$ to indicate that $\omega=:\omega_{X}$ is the Hermitian form.
 In terms of local holomorphic coordinates $(z_{1},\ldots,z_{n})$ on $X$, we may write $\omega=\sqrt{-1}\sum_{i,j=1}^{n}g_{i\overline{j}}dz_{i}\wedge d\overline{z_{j}}$ locally.
@@
 Now, we consider Case (b), i.e., $(N,\omega_{N})$ is a period domain equipped with the canonical invariant metric and $F$ is a horizontal holomorphic map, i.e., $F:U\to N$ is a holomorphic map such that $dF_{x}(T_{x}(M))$ lies inside the horizontal tangent space $T^{h}_{F(x)}(N)\subset T_{F(x)}(N)$ for all $x\in U$.
-Then, $F(U)\subset N$ is an integral manifold for the horizontal distribution on the period domain $N$.
+As in Case (a), we may assume without loss of generality that $p=m$ so that $F^{*}\omega_{N}^{m}=\lambda\omega_{M}^{m}$ holds, and thus $F$ is a holomorphic immersion.
+Then, $F(U)\subset N$ is an integral submanifold for the horizontal distribution on the period domain $N$.
 By a result of Carlson-Toledo [CT88, Proposition 5.3], we have $dF_{x}(T_{x}(M)),dF_{x}(T_{x}(M))]=0$ for $x\in U$.
@@
 for some homogeneous polynomials $G^{j}_{l}$, $1\leq l\leq N^{\prime}_{j}$, of degree $\geq 2$ (cf. Chan-Xiao-Yuan [CXY17, Section 4]). By using Mok [Mo14, Lemma 3 and Lemma 5], we compute
-$\begin{split}\widetilde{\phi}_{m}(\zeta)=&\det\begin{bmatrix}{\lambda_{1}\over(1-|\zeta|^{2})^{2}}&&&&&\\ &{\lambda_{1}\over 1-|\zeta|^{2}}{\bf I}_{p_{1}}&&&&\\ &&\lambda_{1}{\bf I}_{m_{1}-p_{1}-1}&&&\\ &&&\lambda_{2}{\bf I}_{m_{2}}&&\\ &&&&\ddots&\\ &&&&&\lambda_{k}{\bf I}_{m_{k}}\end{bmatrix}\\ =&{\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{p_{1}+2}},\end{split}$
+$\tiny\widetilde{\phi}_{m}(\zeta)=\det\begin{bmatrix}{\lambda_{1}\over(1-|\zeta|^{2})^{2}}&&&&&\\ &{\lambda_{1}\over 1-|\zeta|^{2}}{\bf I}_{p_{1}}&&&&\\ &&\lambda_{1}{\bf I}_{m_{1}-p_{1}-1}&&&\\ &&&\lambda_{2}{\bf I}_{m_{2}}&&\\ &&&&\ddots&\\ &&&&&\lambda_{k}{\bf I}_{m_{k}}\end{bmatrix}={\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{p_{1}+2}},$
 which does not lie in $\Lambda(\Delta)$ by Umehara [Um88], a plain contradiction. Here, $p_{1}$ is a non-negative integer depending only on $D_{1}$. Hence, $\phi_{m}\not\in\Lambda(D)$. ∎